A sample of 21 tree heights gave a sample mean of 26.1 m. Suppose the population standard deviation is 6.2 m. Construct a 80% confidence interval for the population mean tree height. We can assume that the tree heights are normally distributed.

Round to one digit after the decimal point.

Answer: We are 80% confident that the true population mean tree height lies somewhere between      metres and      metres.
Answer #1

\bar{x}\pm Z*\frac{\sigma }{\sqrt{n}}= 26.1 \pm 1.282 *(\frac{ 6.2 }{\sqrt{ 21 }})=( 24.366 , 27.834 )

We are 80% confident that the true population mean tree height lies somewhere between 24.4 metres and 27.8 metres.